OpenStudy (anonymous):

A recipe that makes 4 servings calls for 3/4 cup of apple juice and 2/3 cup of cranberry juice.

OpenStudy (anonymous):

A. 5/7 B. 17/48 C. 1 5/12 D. 5 2/3

OpenStudy (anonymous):

@math&ing001 @mathmath333

OpenStudy (anonymous):

@math&ing001

OpenStudy (anonymous):

C

OpenStudy (anonymous):

can you explain?

OpenStudy (mathmath333):

the question seems incomplete

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

i for got to type ow many cups of juice are there per serving?

OpenStudy (anonymous):

how*

OpenStudy (anonymous):

is that better?

OpenStudy (mathmath333):

rewrite the complete question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A recipe that makes 4 servings calls for 3/4 cup of apple juice and 2/3 cup of cranberry juice. How many cups of juice are there per serving? A. 5/7 B. 17/48 C. 1 5/12 D. 5 2/3

OpenStudy (anonymous):

thats it

OpenStudy (mathmath333):

\(\Large\rm \dfrac{3}{4}a+ \dfrac{2}{3}c=4s \) \(\Large\rm \dfrac{3}{4}\times \dfrac{1}{4}a+ \dfrac{2}{3}\times \dfrac{1}{4}c =4s\times\dfrac{1}{4} \) \(\Large\rm ans= \dfrac{3}{4}\times \dfrac{1}{4}+ \dfrac{2}{3}\times\dfrac{1}{4} \)

OpenStudy (anonymous):

D?

OpenStudy (mathmath333):

wait

OpenStudy (anonymous):

mkai

OpenStudy (anonymous):